# A Linearly Convergent Frank-Wolfe-type Method for Smooth Convex Minimization over the Spectrahedron

### v1

@@ 3.1 Proof of Theorem 2
Using the quadratic growth assumption and the smoothness of $f$ we have that,
$\|{\nabla_{t}-\nabla{}f({\mathbf{X}}^{*})}\|_{F}\leq\beta\|{{\mathbf{X}}_{t}-{\mathbf{X}}^{*}}\|_{F}\leq\frac{\sqrt{\alpha}\beta}{\sqrt{2}}\sqrt{h_{t}}.$
Thus, whenever $h_{t}\leq\frac{\alpha\delta^{2}}{18\beta^{2}}$, we indeed have that $\|{\nabla_{t}-\nabla{}f({\mathbf{X}}^{*})}\|_{F}\leq\frac{\delta}{3}$. Combining this with Eq. (22) and the fact that in case a drop step is not taken we have that $f({\mathbf{X}}_{t+1})\leq f({\mathbf{X}}_{t+1}^{\textrm{FW}})$, we obtain Result (7).
We now move on to prove the main result of the theorem — Result (9). Recall in this case we assume $n>r^{*}\geq 2$. Let us fix some iteration $t$ of Algorithm 1 and denote ${\mathbf{X}}^{*}=\arg\min_{{\mathbf{Y}}\in\mathcal{X}^{*}}\|{{\mathbf{Y}}-{\mathbf{X}}_{t}}\|_{F}$. It holds that,
$\forall i\in[n]:\quad|{\lambda_{i}(\nabla_{t})-\lambda_{i}(\nabla{}f({\mathbf{X}}^{*}))}|$ $\leq\|{\nabla_{t}-\nabla{}f({\mathbf{X}}^{*})}\|\underset{(a)}{\leq}\beta\|{{\mathbf{X}}_{t}-{\mathbf{X}}^{*}}\|_{F}$
$\underset{(b)}{\leq}\frac{\sqrt{2}\beta}{\sqrt{\alpha}}\sqrt{h_{t}},$ (23)
where (a) is due the the smoothness of $f$, and $(b)$ is due to the quadratic growth assumption.
Thus, if $h_{t}\leq\epsilon_{1}:=\frac{\alpha{}\delta^{2}}{32\beta^{2}}$, we have that
$\delta_{t}:=\lambda_{n-r^{*}}(\nabla_{t})-\lambda_{n-r^{*}+1}(\nabla_{t})\geq\delta-2\frac{\sqrt{2}\beta}{\sqrt{\alpha}}\sqrt{\epsilon_{1}}=\frac{\delta}{2}.$ (24)
Denote $\epsilon_{2}=\min\{\frac{(\delta/2)^{2}}{18\beta},\frac{\delta/2}{6}\}$. Using Lemma 3 we have that if $h_{t}<\min\{\epsilon_{1},\epsilon_{2}\}$ and $\textrm{rank}({\mathbf{X}}_{t})>r^{*}$, then a drop step will be preformed. Note also that if $h_{t}<\min\{\epsilon_{1},\epsilon_{2}\}$ and a drop step is not taken, then it must hold that $\textrm{rank}({\mathbf{X}}_{t})\leq r^{*}$. Note also that if $h_{t}<\epsilon_{3}:=\frac{\alpha\lambda_{r^{*}}^{2}}{8}$, using the quadratic growth assumption we have that,
@@
In the complementing case that $\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\perp}{\mathbf{X}}_{t})\leq\frac{h_{t}}{4G}$, it follows from Lemma 6 that
@@
Additionally, using Eq. (3.1) again we have that,
$\lambda_{n-r^{*}+1}(\nabla_{t})-\lambda_{n}(\nabla_{t})$ $\leq\lambda_{n-r^{*}+1}(\nabla{}f({\mathbf{X}}^{*}))-\lambda_{n}(\nabla{}f({\mathbf{X}}^{*}))+\frac{2\sqrt{2}\beta}{\sqrt{\alpha}}\sqrt{h_{t}}$
=$\underset{(a)}{=}\frac{2\sqrt{2}\beta}{\sqrt{\alpha}}\sqrt{h_{t}}\leq\frac{2\sqrt{2}\beta}{\sqrt{\alpha}}\sqrt{\epsilon_{3}}=\beta\lambda_{r^{*}}\underset{(b)}{\leq}2\beta\lambda_{r^{*}}({\mathbf{X}}_{t}),$ (28)
where (a) follows since due to the optimality of ${\mathbf{X}}^{*}$ it holds that $\lambda_{n}(\nabla{}f({\mathbf{X}}^{*}))=\lambda_{n-r^{*}+1}(\nabla{}f({\mathbf{X}}^{*}))$ (see also Lemma 5.2 in [13]), and (b) follows from Eq. (25).
@@
[…]
@@ 4 Numerical Experiments
We conduct numerical experiments with synthetic data for the problem of recovering a ground truth matrix in $\mathcal{S}^{n}$ from random linear measurements. We use a smilier setup and methodology to the one considered in previous works [14, 7]. Concretely, we consider the following optimization problem:
$\min_{{\mathbf{X}}\in\mathcal{S}^{n}}\{f({\mathbf{X}}):=\frac{1}{2}\sum_{i=1}^{m}\left({{\mathbf{a}}_{i}^{\top}(\tau{\mathbf{X}}){\mathbf{a}}_{i}-{\mathbf{b}}_{i}}\right)^{2}\}.$ (30)
@@
We let ${\mathbf{X}}_{\sharp}={\mathbf{U}}_{\sharp}{\mathbf{U}}_{\sharp}^{\top}\in\mathcal{S}^{n}$ denote the ground truth matrix, where ${\mathbf{U}}_{\sharp}\in\mathbb{R}^{n\times r^{*}}$ is first set to a random matrix with standard Gaussian entries and then normalized to have unit Frobenius norm. Each ${\mathbf{a}}_{i}\in\mathbb{R}^{n}$ is a random vector with standard Gaussian entries. We set ${\mathbf{b}}={\mathbf{b}}_{\sharp}+{\mathbf{z}}$, where ${\mathbf{b}}_{\sharp}(i)={\mathbf{a}}_{i}^{\top}{\mathbf{X}}_{\sharp}{\mathbf{a}}_{i}$, and ${\mathbf{z}}=\frac{\|{{\mathbf{b}}_{\sharp}}\|}{2}{\mathbf{v}}$ for a random unit vector ${\mathbf{v}}$.
@@
In order to avoid over-fitting the noise we set the scaling parameter $\tau$ in (30) to $\tau=0.5$ (as also been done in [14, 7]).
@@
We solve Problem (30) to accuracy of the order $10^{-11}$ or lower and denote our estimate for the optimal solution by $\widehat{{\mathbf{X}}^{*}}$ 444this is verified by computing the dual gap for the candidate solution $\widehat{{\mathbf{X}}^{*}}$, which is given by $\langle{\widehat{{\mathbf{X}}^{*}},\nabla{}f(\widehat{{\mathbf{X}}^{*}})}\rangle-\lambda_{n}(\nabla{}f(\widehat{{\mathbf{X}}^{*}}))$ and is always an upper-bound on the approximation error $f(\widehat{{\mathbf{X}}^{*}})-f^{*}$ due the convexity of $f$. We set the smoothness parameter $\beta$ to $\beta=n^{2}/2$ 555while this might seem conservative, note that a simple calculation shows that $\beta\leq\sum_{i=1}^{m}\|{{\mathbf{a}}_{i}}\|^{4}$ and thus we should roughly take $\beta\approx mn^{2}$, however this turns out to be an overly pessimistic estimate (see also experiments in [7]).
Each one of the figures presented below is the average of 10 i.i.d. runs. In particular, we verify that in each of our randomized experiments the strict complementarity condition indeed holds with substantial positive values (see also Section 2.2.1 in [14] for a detailed discussion about verifying the strict complementarity assumption). An exception is Figure 4 which examines a setting in which the strict complementarity condition intentionally does not hold.
@@ 4.1 Comparison to standard Frank-Wolfe
We begin by comparing our Algorithm 1 (ALG 1 in the figures below) to the standard Frank-Wolfe with line-search method (FW in the figures below). Recall that FW has a worst case convergence rate of $O(1/t)$ [20]. In [14] it was established that under strict complementarity and in the special case $r^{*}=1$, it enjoys a linear convergence rate, while it remained unknown if this is also the case for $r^{*}\geq 2$. Figure 1, which plots for each method the value $\log({f({\mathbf{X}}_{t})-f(\widehat{{\mathbf{X}}^{*}})})$ vs the number of iteration $t$, shows that indeed in case $r^{*}=1$ both FW and our Algorithm 1 clearly exhibit a linear convergence rate. However, once we increase $r^{*}$, we see that while our Algorithm 1 maintains a linear convergence rate, FW converges with a sub-linear rate. This suggests that, similarly to the state-of-affairs in the case of optimization over polytopes [12], the standard FW does not exhibit linear convergence once the optimal solution is not an extreme point of the feasible set.
(a) $r^{*}=1$
(b) $r^{*}=2$
(c) $r^{*}=5$
Figure 1: Comparison of Frank-Wolfe with line-search and our Algorithm 1. We set $n=100$ and $m=15nr^{*}$.
@@ 4.2 Comparison of variants of Algorithm 1
We turn to provide empirical motivation for the design choices in our Algorithm 1. In particular, Algorithm 1 prioritises drop steps over other steps (even if these will reduce the objective value slower) and in addition to away steps (which are a classical concept in Frank-Wolfe-based methods [16, 21]), uses also randomized proximal pairwise steps (which also require an estimate for the smoothness parameter $\beta$).
Figure 2 demonstrates how each one of the above mentioned design choices contributes significantly to the fast convergence our Algorithm 1. In particular, we see that the ALG 1 - away variant, which is the natural extension of the Frank-Wolfe with away steps method and is known to converge linearly for polytopes [16, 21, 12], converges only with a sub-linear rate in our experiments.
@@
Figure 2: Comparison of different variants of our Algorithm 1. We set $n=100$, $r^{*}=5$, and $m=15nr^{*}$.
@@ 4.3 Comparison to the Block Frank-Wolfe method
We also compare our method to Block Frank-Wolfe methods, that is, methods that on each iteration perform an update using a rank-$r$ matrix, for some $r\geq r^{*}$, rather than a rank-one update as in the standard Frank-Wolfe [1, 7]. We compare with the method from [1] (Block-FW in Figure 3) which performs updates of the form ${\mathbf{X}}_{t+1}\leftarrow(1-\eta){\mathbf{X}}_{t}+\eta{\mathbf{V}}_{t}$, where ${\mathbf{V}}_{t}\in\mathcal{S}^{n}$ is a rank-$r$ matrix that is obtained by projecting onto $\mathcal{S}_{n}$ the best rank-$r$ approximation (in Frobenius norm) to the matrix ${\mathbf{M}}_{t}:={\mathbf{X}}_{t}-\frac{1}{\eta\beta}\nabla{}f({\mathbf{X}}_{t})$, and can be computed using only the top $r$ components in the eigen-decomposition of ${\mathbf{M}}_{t}$. While this method converges linearly (under quadratic growth), it requires a tight estimation of $r^{*}$ (otherwise it suffers from computationally-expensive high-rank eigen-decompositions), and it requires a good estimate of both $\alpha,\beta$ to set the step size to $\eta=O(\alpha/\beta)$, which is difficult to properly tune in practice. Here we consider a somewhat ideal implementation of this method: I. we set $r=r^{*}$ (i.e., assume exact knowledge of $r^{*}$), and II. we manually tune the step-size parameter $\eta$ for best performance and set $\eta=0.3$.
We can see in Figure 3 (left panel) that when comparing the (log) approximation error vs. number of iterations, as expected, the Block-FW method converges faster than our Algorithm 1 which only uses rank-one matrix computations. However, when we examine the approximation error vs. the number of rank-one updates (recall our implementation of the Block-FW method does a rank-$r^{*}$ update on each iteration, while our Algorithm 1 does at most a rank-two update), see right panel in Figure 3, we see that our Algorithm 1 is in fact faster.
Figure 3: Comparison of the Block-FW method and our Algorithm 1 for $r^{*}=10$. We set $n=150$ and $m=20nr^{*}$.
4.4 What if strict complementarity does not hold?
Finally, we examine what happens when we solve Problem (30) when setting $\tau=\textrm{Tr}({\mathbf{X}}_{\sharp})=1$ and without measurement noise, i.e., ${\mathbf{b}}={\mathbf{b}}_{\sharp}$. In this case, the ground truth matrix ${\mathbf{X}}_{\sharp}$ is an optimal solution of Problem (30) and in particular $\nabla{}f({\mathbf{X}}_{\sharp})=\mathbf{0}$, meaning the strict complementarity condition does not hold. We see in Figure 4 that in this case, our Algorithm 1 maintains its linear convergence, while the standard Frank-Wolfe with line-search method converges with a sub-linear rate. It remains an open question if the quadratic growth condition alone suffices to prove linear convergence for our method, see also Section 5.
Figure 4: Comparison of Frank-Wolfe with line-search and our Algorithm 1 in case strict complementarity does not hold. We set $n=100$, $r^{*}=3$, and $m=10nr^{*}$.

### v2

@@ 3.1 Proof of Theorem 2
Using the smoothness and convexity of $f$ we have that (see for instance Lemma 10 in the appendix),
$\|{\nabla_{t}-\nabla{}f({\mathbf{X}}^{*})}\|_{F}\leq\sqrt{\beta{}h_{t}}.$
Thus, whenever $h_{t}\leq\frac{\delta^{2}}{9\beta}$, we indeed have that $\|{\nabla_{t}-\nabla{}f({\mathbf{X}}^{*})}\|_{F}\leq\frac{\delta}{3}$. Combining this with Eq. (25) and the fact that in case a drop step is not taken we have that $f({\mathbf{X}}_{t+1})\leq f({\mathbf{X}}_{t+1}^{\textrm{FW}})$, we obtain Result (8).
We now move on to prove the main result of the theorem — Result (10). Recall in this case we assume $n>r^{*}\geq 2$. Let us fix some iteration $t$ of Algorithm 1 and denote ${\mathbf{X}}^{*}=\arg\min_{{\mathbf{Y}}\in\mathcal{X}^{*}}\|{{\mathbf{Y}}-{\mathbf{X}}_{t}}\|_{F}$. Using Weyl’s inequality for the eigenvalues, we have that
$\forall i\in[n]:\quad|{\lambda_{i}(\nabla_{t})-\lambda_{i}(\nabla{}f({\mathbf{X}}^{*}))}|$ $\leq\|{\nabla_{t}-\nabla{}f({\mathbf{X}}^{*})}\|\leq\sqrt{\beta{}h_{t}},$ (26)
where the last inequality is due to the smoothness and convexity of $f$ (see Lemma 10 in the appendix).
Thus, if $h_{t}\leq\epsilon_{1}:=\frac{\delta^{2}}{16\beta}$, we have that
$\delta_{t}:=\lambda_{n-r^{*}}(\nabla_{t})-\lambda_{n-r^{*}+1}(\nabla_{t})\geq\delta-2\sqrt{\beta\epsilon_{1}}=\frac{\delta}{2}.$ (27)
Denote $\epsilon_{2}=\min\{\frac{(\delta/2)^{2}}{18\beta},\frac{\delta/2}{6}\}$. Using Lemma 3 we have that if $h_{t}<\min\{\epsilon_{1},\epsilon_{2}\}$ and $\textrm{rank}({\mathbf{X}}_{t})>r^{*}$, then a drop step will be performed. Note also that if $h_{t}<\min\{\epsilon_{1},\epsilon_{2}\}$ and a drop step is not taken, then it must hold that $\textrm{rank}({\mathbf{X}}_{t})\leq r^{*}$. Note also that if $h_{t}<\epsilon_{3}:=\frac{\alpha\lambda_{r^{*}}^{2}}{8}$, then using Weyl’s inequality again and the quadratic growth assumption we have that,
@@
In the complementing case that $\textrm{Tr}({\mathbf{P}}_{t}^{\perp}{\mathbf{X}}_{t})\leq\frac{h_{t}}{4G}$, it follows from Lemma 6 that
@@
Additionally, using Eq. (26) again we have that,
$\lambda_{n-r^{*}+1}(\nabla_{t})-\lambda_{n}(\nabla_{t})$ $\leq\lambda_{n-r^{*}+1}(\nabla{}f({\mathbf{X}}^{*}))-\lambda_{n}(\nabla{}f({\mathbf{X}}^{*}))+2\sqrt{\beta{}h_{t}}$
$\underset{(a)}{=}2\sqrt{\beta{}h_{t}}\leq 2\sqrt{\beta{}\epsilon_{3}}=\sqrt{\frac{\alpha\beta}{2}}\lambda_{r^{*}}\underset{(b)}{\leq}\sqrt{2}\beta\lambda_{r^{*}}({\mathbf{X}}_{t}),$ (31)
where (a) follows since due to the optimality of $\mathbf{X}$$^{*}$ it holds that $\lambda_{n}(\nabla{}f({\mathbf{X}}^{*}))=\lambda_{n-r^{*}+1}(\nabla{}f({\mathbf{X}}^{*}))$ (see also Lemma 5.2 in [14]), and (b) follows from Eq. (28) and the fact that $\alpha\leq\beta$ (e.g., Lemma 11 in the appendix).
@@
Finally, it remains to prove Result (9) for the case $r^{*}=n$, which follows as a simplified case of Result (10). If $h_{t}<\epsilon_{3}$, then we have that Eq. (28) and Eq. (3.1) hold, and as a consequence the assumptions of Lemma 6 and Lemma 7 hold. Thus, it follows that Eq. (32) also holds.
@@ 4 Numerical Experiments
We conduct numerical experiments with synthetic data for the problem of recovering a low-rank ground truth matrix in $\mathcal{S}^{n}$ from random linear measurements under two noise models: one in which the measurements are perturbed with isotropic (Gaussian) noise, and we use a standard least squares (LS) estimator. We refer to this as the LS setting. In the second model, only a constant fraction of measurements are grossly corrupted, and we use the Huber loss for robust estimation. We refer to this as the Huber setting. We use a similar setup and methodology to the one considered in previous works [15, 8]. Concretely, we consider the following optimization problem:
$\min_{{\mathbf{X}}\in\mathcal{S}^{n}}\{f({\mathbf{X}}):=\sum_{i=1}^{m}\ell\left({{\mathbf{a}}_{i}^{\top}(\tau{\mathbf{X}}){\mathbf{a}}_{i}-{\mathbf{b}}_{i}}\right)\},$ (33)
where in the LS setting $\ell$ is simply the square loss: $\ell(x)=\frac{1}{2}x^{2}$, and in the Huber setting $\ell(x)=H_{\zeta}(x)$ — the Huber loss with parameter $\zeta$777formally defined as: $H_{\zeta}(x)=\left\{\begin{array}[]{ll}\frac{1}{2}x^{2}&\mbox{if $|x|\leq\zeta$};\\ \zeta(|x|-\frac{1}{2}\zeta)&\mbox{if $|x|>\zeta$}.\end{array}\right$..
@@
We let ${\mathbf{X}}_{\sharp}={\mathbf{U}}_{\sharp}{\mathbf{U}}_{\sharp}^{\top}\in\mathcal{S}^{n}$ denote the ground truth matrix, where ${\mathbf{U}}_{\sharp}\in\mathbb{R}^{n\times r^{*}}$ is first set to a random matrix with standard Gaussian entries and then normalized to have unit Frobenius norm. Each ${\mathbf{a}}_{i}\in\mathbb{R}^{n}$ is a random vector with standard Gaussian entries. We set ${\mathbf{b}}={\mathbf{b}}_{\sharp}+{\mathbf{z}}$, where ${\mathbf{b}}_{\sharp}(i)={\mathbf{a}}_{i}^{\top}{\mathbf{X}}_{\sharp}{\mathbf{a}}_{i}$ and $\mathbf{z}$} is additive noise. In the LS setting we set ${\mathbf{z}}=\frac{\|{{\mathbf{b}}_{\sharp}}\|}{2}{\mathbf{v}}$ for a random unit vector $\mathbf{v}$}. In the Huber setting, for each $i\in[m]$ we set with probability $0.15$, ${\mathbf{z}}(i)=\pm 4.2\frac{\|{{\mathbf{b}}_{\sharp}}\|}{\sqrt{m}}$, and ${\mathbf{z}}(i)=0$ with remaining probability. We set the Huber loss parameter to $\zeta=0.7\frac{\|{{\mathbf{b}}_{\sharp}}\|}{\sqrt{m}}$.
@@
In order to avoid over-fitting the noise, in both the LS and Huber settings, we set the scaling parameter $\tau$ in (33) to $\tau=0.5$ (as also been done in [15, 8]).
@@
We solve Problem (33) to accuracy of the order $10^{-10}$ or lower and denote our estimate for the optimal solution by $\widehat{{\mathbf{X}}^{*}}$ 888this is verified by computing the dual gap for the candidate solution $\widehat{{\mathbf{X}}^{*}}$, which is given by $\langle{\widehat{{\mathbf{X}}^{*}},\nabla{}f(\widehat{{\mathbf{X}}^{*}})}\rangle-\lambda_{n}(\nabla{}f(\widehat{{\mathbf{X}}^{*}}))$ and is always an upper-bound on the approximation error $f(\widehat{{\mathbf{X}}^{*}})-f^{*}$ due to the convexity of $f$. For both the LS and Huber settings we verify that in each of our randomized experiments the strict complementarity condition indeed holds with substantial positive values (see also Section 2.2.1 in [15] for a detailed discussion about verifying the strict complementarity assumption).
We also consider a third setting, which we refer to as the No SC setting, in which strict complementarity does not hold. In this setting we do not add noise to the measurements (i.e., we set ${\mathbf{b}}={\mathbf{b}}_{\sharp}$) and we also set $\tau=1$. This guarantees that at the optimal solution to (33), the gradient is zero, and in particular strict complementarity cannot hold.
We set the smoothness parameter $\beta$ to $\beta=n^{2}/2$ 999while this might seem conservative, note that a simple calculation shows that $\beta\leq\sum_{i=1}^{m}\|{{\mathbf{a}}_{i}}\|^{4}$ and thus we should roughly take $\beta\approx mn^{2}$, however this turns out to be an overly pessimistic estimate (see also experiments in [8]).
Each one of the figures presented below is the average of 10 i.i.d. runs, where on each run we resample both the data to Problem (33) and the randomness in our Algorithm 1 .
@@ 4.1 Comparison to standard Frank-Wolfe
We begin by comparing our Algorithm 1 (ALG 1 in the figures below) to the standard Frank-Wolfe with line-search method (FW in the figures below). Recall that FW has a worst case convergence rate of $O(1/t)$ [19]. In [15] it was established that under strict complementarity and in the special case $r^{*}=1$, it enjoys a linear convergence rate, while it remained unknown if this is also the case for $r^{*}\geq 2$. Figure 1, which plots for each method the value $\log({f({\mathbf{X}}_{t})-f(\widehat{{\mathbf{X}}^{*}})})$ vs the number of iterations $t$, shows that indeed in case $r^{*}=1$, in both the LS and Huber settings, both FW and our Algorithm 1 clearly exhibit a linear convergence rate. However, once we increase $r^{*}$, we see that while our Algorithm 1 maintains a linear convergence rate, FW converges only with a sub-linear rate. Additionally, in the No SC setting, we see that even for $r^{*}=1$, FW converges sub-linearly, while our algorithm exhibits a linear convergence rate. This suggests that, similarly to the state-of-affairs in the case of optimization over polytopes [13], the standard FW does not exhibit linear convergence once the optimal solution is not an extreme point of the feasible set.
(a) LS, $r^{*}=1$
(b) LS, $r^{*}=2$
(c) LS, $r^{*}=5$
(d) Huber, $r^{*}=1$
(e) Huber, $r^{*}=2$
(f) Huber, $r^{*}=5$
(g) No SC, $r^{*}=1$
(h) No SC, $r^{*}=2$
(i) No SC, $r^{*}=5$
Figure 1: Comparison of Frank-Wolfe with line-search and our Algorithm 1. We set $n=100$ and $m=15nr^{*}$ in all experiments.
@@ 4.2 Comparison of variants of Algorithm 1
We turn to provide empirical motivation for the design choices in our Algorithm 1. In particular, Algorithm 1 prioritizes drop steps over other steps (even if these will reduce the objective value more slowly) and in addition to away steps (which are a classical concept in Frank-Wolfe-based methods [17, 22], though recall our away steps take a different form than those for polytopes), uses also randomized proximal pairwise steps (which also require an estimate for the smoothness parameter $\beta$).
Figure 2 demonstrates how each one of the above mentioned design choices contributes to the fast convergence of our Algorithm 1. In particular, we see that the ALG 1 - away variant, which is a natural extension of the Frank-Wolfe with away steps method that is known to converge linearly for polytopes [17, 22, 13], converges only with a sub-linear rate in our experiments. Interestingly, the ALG 1 - noaway variant, which does not use away steps at all (but does use drop steps), performs comparably to Algorithm 1 in both the LS and Huber settings, however, it performs considerably worse in the No SC setting (when $r^{*}>2$).
@@
(a) LS, $r^{*}=2$
(b) LS, $r^{*}=5$
(c) LS, $r^{*}=7$
(d) Huber, $r^{*}=2$
(e) Huber, $r^{*}=5$
(f) Huber, $r^{*}=8$
(g) No SC, $r^{*}=2$
(h) No SC, $r^{*}=5$
(i) No SC, $r^{*}=8$
Figure 2: Comparison of different variants of our Algorithm 1. In all experiments, except for Huber $r^{*}=8$, we set $n=100$, $m=15nr^{*}$. In Huber $r^{*}=8$ we set $n=100$, $m=20nr^{*}$.
@@ 4.3 Comparison to the Block Frank-Wolfe method
We also compare our method to Block Frank-Wolfe methods, that is, methods that on each iteration perform an update using a rank-$r$ matrix, for some $r\geq r^{*}$, rather than a rank-one update as in the standard Frank-Wolfe [1, 8]. We compare with the method from [1] (Block-FW in Figure 3) which performs updates of the form ${\mathbf{X}}_{t+1}\leftarrow(1-\eta){\mathbf{X}}_{t}+\eta{\mathbf{V}}_{t}$, where ${\mathbf{V}}_{t}\in\mathcal{S}^{n}$ is a rank-$r$ matrix that is obtained by projecting onto $\mathcal{S}^{n}$ the best rank-$r$ approximation (in Frobenius norm) to the matrix ${\mathbf{M}}_{t}:={\mathbf{X}}_{t}-\frac{1}{\eta\beta}\nabla{}f({\mathbf{X}}_{t})$, and can be computed using only the top $r$ components in the eigen-decomposition of ${\mathbf{M}}_{t}$. While this method converges linearly (under quadratic growth), it requires a tight estimation of $r^{*}$ (otherwise it suffers from computationally-expensive high-rank eigen-decompositions), and it requires a good estimate of both $\alpha,\beta$ to set the step size to $\eta=O(\alpha/\beta)$, which is difficult to properly tune in practice. Here we consider a somewhat ideal implementation of this method: I. we set $r=r^{*}$ (i.e., assume exact knowledge of $r^{*}$), and II. we manually tune the step-size parameter $\eta$ for best performance and set $\eta=0.3$.
We can see in Figure 3 (left panels) that when comparing the (log) approximation error vs. number of iterations, as expected, the Block-FW method converges faster than our Algorithm 1 which only uses rank-one matrix computations. However, when we examine the approximation error vs. the number of rank-one updates (recall our implementation of the Block-FW method does a rank-$r^{*}$ update on each iteration, while our Algorithm 1 does at most a rank-two update), see right panels in Figure 3, we see that our Algorithm 1 is in fact faster.
(a) LS, $r^{*}=5$
(b) LS, $r^{*}=5$
(c) LS, $r^{*}=7$
(d) LS, $r^{*}=7$
(e) Huber, $r^{*}=5$
(f) Huber, $r^{*}=5$
(g) Huber, $r^{*}=8$
(h) Huber, $r^{*}=8$
Figure 3: Comparison of the Block-FW method and our Algorithm 1. In all experiments, except for Huber $r^{*}=8$, we set $n=100$, $m=15nr^{*}$. In Huber $r^{*}=8$ we set $n=100$, $m=20nr^{*}$.